Convert 3 yards to a) meter and b) mm


1
Expert's answer
2021-03-31T09:39:54-0400

(a)

"3\ yard\cdot\dfrac{0.9144\ m}{1\ yard}=2.7432\ m."

(b)

"3\ yard\cdot\dfrac{0.9144\ m}{1\ yard}\cdot\dfrac{1000\ mm}{1\ m}=2743.2\ mm."